OpenStudy (anonymous):

solve for L \[d=\frac{ LR^2 }{ R^2+R^1 } \]

OpenStudy (abmon98):

get rid of thedenominator by multiplying both sides d(R^2+R^1)=LR^2 Divide by R^2 d(R^2+R1)/R^2=L

OpenStudy (akashdeepdeb):

\[d = \frac{LR^2}{R^2 + R}\] Cross multiply \[d(R^2 + R) = LR^2\] Divide both sides by \(R^2\) to get rid of the \(R^2\) on the right hand side, to get the value of \(L\). \[\frac{d(R^2+R)}{R^2} = L\] \[L = d(1 + \frac{1}{R})\] Understood this? :)

OpenStudy (akashdeepdeb):

Is that \(R^1\) or \(R_1\) ? Manipulate accordingly. The steps are the same. :)

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

so whats next?

OpenStudy (akashdeepdeb):

Is it \(R^2\) or \(R_2\) ?

OpenStudy (akashdeepdeb):

I am guessing it is the subscript. So, \(R_2, R_1, d, L\) are completely different terms. So, we have to solve for: \[d = \frac{LR_2}{R_2 + R_1}\] Now, to find 'L' how can we isolate it and represent it in terms of the other variables? Step 1: Well, we can cross multiple right? [Same as multiplying \(R_2 + R_1\)on both sides] \[d(R_2 + R_1) = \frac{LR_2}{(R_2 + R_1)} (R_2 + R_1)\] \[d(R_2 + R_1) = \frac{LR_2}{\cancel{(R_2 + R_1)}} \cancel{(R_2 + R_1)}\] \[d(R_2 + R_1) = LR_2\] Step 2: I see L there, but it is multiplied to \(R_2\) [Bahh!! -_-] How can I get rid of the \(R_2\) ? Well, for one, dividing has the opposite effect of multiplying! Let's try that! \[d \frac{(R_2 + R_1)}{R_2} = L\] Step 3: Not necessary, but can be used to make it look more mathematical :-) \[L = d (1 + \frac{R_1}{R_2}) \] Getting this? :)